# Surrogate-Assisted Targeted Learning for Delayed Outcomes under Administrative Censoring

### 11 Mar 2026 v1

@@ 4.1 Stage 1: Initial Estimation via Super Learner
Remark 6 (Observed-Data Regression vs. Counterfactual Evaluation of $\bar{Q}_{\mathrm{int}}$).
Equation (31) defines $\bar{Q}_{\mathrm{int}}(A,W,t)$ as an observed-data regression of $\bar{Q}_{Y}$ on the observed treatment $A$. During estimation, this regression is fit on the full data using the observed $A\in\{0,1\}$. Counterfactual predictions are then obtained by evaluating the fitted model at the fixed intervention values $A=1$ and $A=0$ separately, yielding $\hat{\bar{Q}}_{\mathrm{int}}(1,W,t)$ and $\hat{\bar{Q}}_{\mathrm{int}}(0,W,t)$. These counterfactual evaluations enter the ATE estimator (37). The regression is fit on the observed scale and evaluated on the interventional scale; conflating these two steps would produce a biased plug-in estimator, which is precisely what the TMLE fluctuation step in Section 4.2 corrects.
As established in Remark 4, the treatment propensity $g_{A}(1\mid W,t)=(t-1)/(T-1)$ is known from the randomization schedule and is plugged in directly. Only $g_{\Delta}$ requires data-adaptive estimation.
@@ 5.2 Asymptotic Normality
Asymptotic linearity (43) follows from the von Mises expansion of $\hat{\Psi}_{\mathrm{TMLE}}$ around $P_{0}$. The TMLE fluctuation step is constructed precisely to solve the efficient score equation $J^{-1}\sum_{j}\mathrm{EIC}_{j}=0$ (up to the convergence tolerance (36)), which eliminates the first-order bias term. The residual is the second-order remainder $R_{2}(\hat{P},P_{0})$ characterized in Supplement, Section S2.4. Under condition (C3), $R_{2}=o_{P}(J^{-1/2})$; this follows from the bilinear form of $R_{2}$ established in Supplement, Section S2.4 and the Cauchy–Schwarz inequality applied to each product of $L^{2}(P_{0})$ errors. When the CV-TMLE of Supplement, Section S3 is used, condition (C3) is achieved without requiring a Donsker condition on the nuisance function classes, because sample splitting eliminates the empirical process terms that would otherwise arise (Zheng and van der Laan, 2011). Asymptotic normality (44) and consistency of the sandwich variance estimator then follow from the Lindeberg–Feller CLT applied to the i.i.d. cluster scores $\{\mathrm{EIC}_{j}\}_{j=1}^{J}$; the complete argument is given in Supplement, Section S2.4.∎∎
Remark 7 (Double robustness of the asymptotic result).
The product-rate condition (C3) is doubly robust in the following sense (Bang and Robins, 2005; Tsiatis, 2006): it is satisfied if either (i) the outcome models $(\bar{Q}_{Y},\bar{Q}_{\mathrm{int}})$ are consistently estimated at any rate, or (ii) the propensity mechanisms $(g_{A},g_{\Delta})$ are consistently estimated at any rate. In particular, since $g_{A}$ is known by design in the SW-CRT (Remark 4), the condition reduces to requiring only that $\hat{\bar{Q}_{Y}}$ and $\hat{\bar{Q}_{\mathrm{int}}}$ converge at rate $o_{P}(J^{-1/4})$, a strictly weaker requirement than pointwise consistency.
Proposition 2 (Nested Cross-Product Term in the DML Remainder).
Remark 8 (Why DML cross-fitting does not eliminate $R_{SY}$).
[…]
@@ 5.3 Non-asymptotic Coverage Bound
Beyond asymptotic validity, the bridge-based estimator admits an explicit non-asymptotic coverage analysis at the cluster level. The resulting Berry–Esseen bound (Theorem S1, Supplementary Material) quantifies how bounded cluster scores, variance concentration, and nuisance-estimation remainder jointly determine finite-sample undercoverage, thereby linking the semiparametric theory directly to the small-$J$ regimes encountered in practice. Corollary S1 gives the minimum cluster count $J^{*}$ for a target coverage level; for the simulation DGP of Section 6, $J^{*}\approx 27$ for 1% maximum undercoverage. Corollary S2 (Supplement, Section S2.7) extends this to characterize how nuisance convergence rates govern the sandwich variance gap, showing that the constant coverage shortfall observed in Block I simulations is the finite-sample signature of estimation near the product-rate boundary.
@@ 6 Monte Carlo Simulation Study
The simulation study is designed to examine the main qualitative implications of the theory: relative stability of the surrogate-bridge estimator as observation probabilities become small (Blocks I and III), the finite-sample behavior of the asymptotic double-robustness theory under deliberate nuisance misspecification (Block II), and the relationship between coverage and the remainder-variance analysis of Section 5.3. The three scenario blocks are organized around increasing administrative censoring severity, misspecified nuisance models, and varying cluster counts. Throughout, “double robustness” refers to the asymptotic identification of the first-order remainder; finite-sample performance can still depend strongly on which nuisance component is misspecified and on the magnitude of the residual second-order error.
@@ 6.1 Data-Generating Process
Trial design.
We simulate a standard SW-CRT with $J$ clusters, $T=7$ calendar time steps, and $n_{j}=40$ individuals per cluster per step, giving a total sample size of $N=J\times T\times n_{j}$. Crossover times $\tau_{j}$ are drawn uniformly without replacement from $\{2,3,4,5,6,7\}$, so that exactly J/(T1)J/(T-1) clusters (rounded to the nearest integer) cross over at each step. All clusters begin in the control condition: $A_{jt}=\mathbf{1}(t\geq\tau_{j})$. Administrative censoring of the delayed outcome is structurally imposed: an individual in cluster $j$ at time step $t$ can contribute an observed delayed outcome only if $\tau_{j}+t_{\mathrm{lag}}\leq T+t_{\mathrm{grace}}$, where $t_{\mathrm{lag}}=2$ steps is the outcome delay and $t_{\mathrm{grace}}=1$ step is an administrative grace period. This implies that clusters crossing over at steps T1T-1 and $T$ have near-zero probability of observing $Y$, creating the design-induced near-boundary observation regime described in Section 1.
Cluster random effect.
The cluster-level random effect is drawn independently as $b_{j}\overset{\mathrm{i.i.d.}}{\sim}\mathcal{N}(0,\sigma_{b}^{2})$ with $\sigma_{b}=0.184$ ($\sigma_{b}^{2}=0.034$, $\mathrm{ICC}\approx 0.05$), consistent with the ICC model of Section 2.3 and empirical ICC levels typical of public-health SW-CRTs (Hemming and Taljaard, 2020).
Baseline covariates.
Each individual has a $p=3$ dimensional baseline covariate vector:
$W_{ij}=(W_{ij,1},\,W_{ij,2},\,W_{ij,3})$
where $W_{ij,1}\sim\mathcal{N}(0,1)$, $W_{ij,2}\sim\mathrm{Bernoulli}(0.4)$, and $W_{ij,3}\sim\mathrm{Uniform}(0,1)$, all independent across individuals and clusters.
Secular time trend.
We specify the true secular time trend as a non-linear function of the calendar step to represent the realistic scenario where a linear time trend is misspecified:
$\lambda(t)=0.5\sin\!\left(\frac{\pi(t-1)}{T-1}\right)+0.3\left(\frac{t-1}{T-1}\right)^{2}.$
This ensures that parametric estimators using a simple linear-in-$t$ fixed effect will exhibit systematic secular time confounding bias.
Short-term surrogate.
The surrogate is generated from a linear model with a cluster random effect, treatment effect, covariate effects, and secular time trend:
$S_{ijt}$ $=0.8\,A_{jt}+0.4\,W_{ij,1}-0.3\,W_{ij,2}+0.5\,\lambda(t)+0.6\,b_{j}+\varepsilon^{S}_{ijt},$
$\varepsilon^{S}_{ijt}$ $\overset{\mathrm{i.i.d.}}{\sim}\mathcal{N}(0,\,0.5^{2}).$ (48)
The coefficient $0.6$ on $b_{j}$ induces a strong within-cluster surrogate–outcome correlation, making $S$ a useful bridge for the missing delayed outcome.
Observation indicator.
Administrative censoring follows a logistic model consistent with Assumption 3:
$\operatorname{logit}\bigl(\mathbb{P}(\Delta_{ijt}=1\mid S_{ijt},A_{jt},W_{ij},t)\bigr)$ $=3.0-2.5\cdot\mathbf{1}(t\geq T-1)-0.8\cdot\mathbf{1}(t=T)$
$\quad+0.4\,S_{ijt}+0.3\,A_{jt}-0.2\,W_{ij,3}.$ (49)
The large negative late-step coefficients concentrate censoring in the final two steps (overall rate $\approx 28\%$, rising to $\approx 80\%$ at $t=T$), creating the near-boundary observation regime for late-crossing clusters.
Delayed primary outcome.
Following the structural equation (6) and the ICC decomposition of Section 2.3:
$Y_{ijt}$ $=\Psi_{0}\,A_{jt}+0.5\,S_{ijt}+0.4\,W_{ij,1}-0.2\,W_{ij,2}+0.6\,W_{ij,3}+0.8\,\lambda(t)+b_{j}+\varepsilon^{Y}_{ijt},$
$\varepsilon^{Y}_{ijt}$ $\overset{\mathrm{i.i.d.}}{\sim}\mathcal{N}(0,\,0.8^{2}),$ (50)
where $\Psi_{0}=-0.28$ is the direct-effect coefficient, $\mathrm{ICC}=0.034/(0.034+0.64)\approx 0.05$.
True ATE.
The true total causal ATE is $\Psi^{*}=-0.28+0.5\times 0.8=0.12$ (direct effect plus mediated path through $S$), placed near the detection boundary for $J\approx 30$.
@@ 6.2 Simulation Scenarios
$J=30$, $\Psi^{*}=0.12$. We consider three misspecification scenarios designed to probe the double robustness property of Theorem 2:
  1. (i)
    Outcome model misspecified, propensities correctly specified. The outcome models $\bar{Q}_{Y}$ and $\bar{Q}_{\mathrm{int}}$ are fit using a main-effects linear regression that omits the non-linear time trend $\lambda(t)$ and the interaction $S_{ijt}\times A_{jt}$; $g_{\Delta}$ is estimated by the correctly specified logistic model. Under the asymptotic double-robustness theory, consistency is still possible when the propensity side is correctly specified, but finite-sample performance may deteriorate if the misspecified outcome model leaves a large structured remainder.
  2. (ii)
    Propensities misspecified, outcome models correctly specified. $g_{\Delta}$ is estimated by logistic regression omitting $S_{ijt}$ (violating the surrogate-conditional independence structure of Assumption 3); $\bar{Q}_{Y}$ and $\bar{Q}_{\mathrm{int}}$ are correctly specified. Because $g_{A}$ is design-based and used as a plug-in (Remark 3), the treatment propensity remains correctly specified; only $g_{\Delta}$ is misspecified. Under the asymptotic double-robustness theory, the estimator is expected to remain stable when the outcome regressions are correctly specified, even if $g_{\Delta}$ is misspecified.
  3. (iii)
    Both nuisance models misspecified. Both the outcome models and $g_{\Delta}$ are misspecified as in (i) and (ii). This scenario is expected to produce estimator bias and is included as a negative control to verify that double robustness does not extend to simultaneous misspecification of both sides.
@@ 6.3 Results
Table 3 reports performance metrics for all three estimators across $J\in\{10,20,30,50,100\}$. The CV-TMLE achieves near-zero bias (<0.004<0.004) at all cluster counts. The GLMM exhibits persistent bias of approximately +0.07+0.07 at $J=30$ due to the misspecified secular time trend, and its coverage collapses from 0.52 at $J=10$ to 0.07 at $J=100$ as the standard error shrinks around the biased point estimate. IPCW exhibits substantial positive bias (approximately +0.22+0.22 at $J=30$) and markedly higher variance than the CV-TMLE, consistent with the theoretical prediction that inverse weighting becomes unstable when $g_{\Delta}$ concentrates near the positivity boundary for late-crossing clusters.
Coverage for CV-TMLE stabilises near 0.87–0.91 across $J\geq 20$, with no systematic improvement as $J$ grows to 100. Table 5 decomposes this gap: the sandwich estimates $\mathrm{Var}(\mathrm{EIC}_{j})/J$ but does not capture the variance of the second-order remainder $R_{2}$. The empirical variance exceeds the mean sandwich by a factor of about 1.6 at all $J$; this magnitude is broadly consistent with the observed coverage shortfall in the 0.87–0.89 range (see predicted column). This ratio is approximately constant because both terms are $O(J^{-1})$ under parametric nuisance rates. The theory suggests that this gap should diminish once nuisance estimation is sufficiently faster than the product-rate boundary; the present simulations indicate that the parametric nuisance specifications used here remain in a regime where that improvement is not yet visible. CV-TMLE power scales from 0.41 at $J=10$ to 1.00 at $J=100$, reflecting strong signal recovery together with very small empirical bias.
@@
Table 2: Variance decomposition for CV-TMLE coverage (Block I). $\widehat{\mathrm{Var}}_{\mathrm{emp}}$: empirical variance of $\hat{\Psi}$ across 1,000 replicates. $\overline{\mathrm{Sand.}}$: mean sandwich variance estimate. Ratio >1>1 indicates sandwich underestimation attributable to unaccounted $R_{2}$ variance. Predicted coverage computed as $P(|Z|<t_{J-1}/\sqrt{\mathrm{Ratio}})$; agreement with actual coverage is approximate.
$J$ $\widehat{\mathrm{Var}}_{\mathrm{emp}}$ ($\times 10^{-3}$) $\overline{\mathrm{Sand.}}$ ($\times 10^{-3}$) Ratio Predicted Cov. Actual Cov.
10 8.44 4.80 1.76 0.912 0.878
20 4.83 2.83 1.71 0.891 0.869
30 2.93 1.84 1.59 0.895 0.892
50 1.80 1.11 1.62 0.886 0.866
100 0.87 0.54 1.61 0.883 0.878
Table 3: Block I simulation results: performance across cluster counts ($\Psi^{*}=0.12$, $T=7$, $n_{j}=40$, 1,000 replicates). “Cov.” denotes 95% CI coverage ($t_{J-1}$ critical values).
$J$ Estimator Bias Var $(\times 10^{-2})$ RMSE Cov. Power
10 GLMM +0.121+0.121 0.41 0.137 0.519 0.970
IPCW +0.235+0.235 0.61 0.248 0.985 0.172
CV-TMLE +0.004+0.004 0.84 0.092 0.878 0.410
20 GLMM +0.104+0.104 0.21 0.114 0.373 0.998
IPCW +0.231+0.231 0.33 0.238 0.934 0.649
CV-TMLE +0.001+0.001 0.48 0.070 0.869 0.554
30 GLMM +0.066+0.066 0.16 0.077 0.597 0.999
IPCW +0.219+0.219 0.22 0.224 0.787 0.939
CV-TMLE +0.000+0.000 0.29 0.054 0.892 0.727
50 GLMM +0.083+0.083 0.08 0.088 0.203 1.000
IPCW +0.224+0.224 0.12 0.227 0.220 1.000
CV-TMLE +0.003+0.003 0.18 0.043 0.866 0.897
100 GLMM +0.074+0.074 0.05 0.077 0.068 1.000
IPCW +0.220+0.220 0.06 0.221 0.000 1.000
CV-TMLE +0.000+0.000 0.09 0.030 0.878 0.997
@@
Additional simulations varying $\sigma_{b}^{2}\in\{0.006,0.034,0.071,0.160\}$ (ICC $\approx$ 0.01–0.20) confirm that the cluster-robust sandwich estimator maintains valid coverage across the empirically observed ICC range in implementation-science SW-CRTs; full results are in Supplement, Section S4.
Table 5: Variance decomposition for CV-TMLE coverage (Block I). $\widehat{\mathrm{Var}}_{\mathrm{emp}}$: empirical variance of $\hat{\Psi}$ across 1,000 replicates. $\overline{\mathrm{Sand.}}$: mean sandwich variance estimate. Ratio >1>1 indicates sandwich underestimation attributable to unaccounted $R_{2}$ variance. Predicted coverage computed as $P(|Z|<t_{J-1}/\sqrt{\mathrm{Ratio}})$; agreement with actual coverage is approximate.
$J$ $\widehat{\mathrm{Var}}_{\mathrm{emp}}$ ($\times 10^{-3}$) $\overline{\mathrm{Sand.}}$ ($\times 10^{-3}$) Ratio Predicted Cov. Actual Cov.
10 8.44 4.80 1.76 0.912 0.878
20 4.83 2.83 1.71 0.891 0.869
30 2.93 1.84 1.59 0.895 0.892
50 1.80 1.11 1.62 0.886 0.866
100 0.87 0.54 1.61 0.883 0.878

### 15 Mar 2026 v2

@@ 4.1 Stage 1: Initial Estimation via Super Learner
$\bar{Q}_{\mathrm{int}}$ is fit on the observed $A$ and evaluated counterfactually at $A=1$ and $A=0$ to produce $\hat{\bar{Q}}_{\mathrm{int}}(1,W,t)$ and $\hat{\bar{Q}}_{\mathrm{int}}(0,W,t)$; the TMLE fluctuation step corrects the resulting plug-in bias. The treatment propensity $g_{A}(1\mid W,t)=(t-1)/(T-1)$ is plugged in directly from the randomization schedule; only $g_{\Delta}$ requires data-adaptive estimation.
@@ 5.2 Asymptotic Normality
The von Mises expansion of $\hat{\Psi}_{\mathrm{TMLE}}$ around $P_{0}$ yields
$\hat{\Psi}_{\mathrm{TMLE}}-\Psi(P_{0})\;=\;\frac{1}{J}\sum_{j=1}^{J}\mathrm{EIC}_{j}\;+\;R_{2}(\hat{P},P_{0})\;+\;R_{SY}(\hat{P},P_{0})\;+\;o_{P}(J^{-1/2}),$ (38)
where $R_{2}$ is the standard bilinear nuisance-error remainder and $R_{SY}$ is the nested cross-product term of Proposition 1.
Elimination of $R_{SY}$. The second fluctuation step (Section 4.2) enforces the score equation $N^{-1}\sum_{j,i,t}D_{S,a}(O_{ijt})=0$ for each $a$. Expanding $D_{S,a}$ from (11) shows that this constraint sets the propensity-weighted empirical mean of $\bar{Q}^{\,*}_{Y}$ equal to $\bar{Q}^{\,*}_{\mathrm{int}}$ at the observed surrogate values, i.e.,
$\frac{1}{N}\sum_{j,i,t}\frac{I(A=a)}{\hat{g}_{A}(a\mid W,t)}\bigl[\bar{Q}^{\,*}_{Y}(S_{ijt},a,W_{ijt},t)-\bar{Q}^{\,*}_{\mathrm{int}}(a,W_{ijt},t)\bigr]\;=\;0.$ (39)
Because $\bar{Q}^{\,*}_{\mathrm{int}}$ is thereby defined as a propensity-weighted empirical average of $\bar{Q}^{\,*}_{Y}$ over the observed surrogate distribution—rather than as an explicit integral against an estimated $\hat{f}_{S}$—the cross-product $(\hat{Q}_{Y}-Q^{0}_{Y})(\hat{f}_{S}-f^{0}_{S})$ that constitutes $R_{SY}$ never enters the expansion; $R_{SY}$ is absorbed into the efficient score without any requirement on the rate of $\hat{f}_{S}$. The formal calculation is in Supplement, Section S2.6.
Control of $R_{2}$. The residual remainder $R_{2}(\hat{P},P_{0})$ is bilinear in the nuisance errors $(\hat{Q}_{Y}-Q^{0}_{Y})$ and $(\hat{g}_{\Delta}-g^{0}_{\Delta})$, and in $(\hat{Q}_{\mathrm{int}}-Q^{0}_{\mathrm{int}})$ and $(\hat{g}_{A}-g^{0}_{A})$. Cauchy–Schwarz gives $|R_{2}|\leq\|\hat{Q}_{Y}-Q^{0}_{Y}\|_{P_{0}}\|\hat{g}_{\Delta}-g^{0}_{\Delta}\|_{P_{0}}+\|\hat{Q}_{\mathrm{int}}-Q^{0}_{\mathrm{int}}\|_{P_{0}}\|\hat{g}_{A}-g^{0}_{A}\|_{P_{0}}$, which is $o_{P}(J^{-1/2})$ by condition (C3). When the CV-TMLE of Supplement, Section S3 is used, sample splitting eliminates the empirical process terms that would otherwise require a Donsker condition (Zheng and van der Laan, 2011).
Asymptotic normality. With $R_{2}+R_{SY}=o_{P}(J^{-1/2})$ established, the Lindeberg–Feller CLT applied to the i.i.d. cluster scores $\{\mathrm{EIC}_{j}\}_{j=1}^{J}$ gives (37). Consistency of the sandwich estimator (32) for $\sigma^{2}/J$ follows by the law of large numbers applied at the cluster level. The complete argument is in Supplement, Section S2.4.∎∎
Remark 3 (Double robustness and role of $R_{SY}$).
Condition (C3) is doubly robust (Bang and Robins, 2005): it holds if either the outcome models $(\bar{Q}_{Y},\bar{Q}_{\mathrm{int}})$ or the propensities $(g_{A},g_{\Delta})$ are consistently estimated. Since $g_{A}$ is known by design, (C3) reduces to $\|\hat{\bar{Q}}_{Y}-\bar{Q}^{0}_{Y}\|_{P_{0}}\cdot\|\hat{g}_{\Delta}-g^{0}_{\Delta}\|_{P_{0}}=o_{P}(J^{-1/2})$. The key separation from DML is (39): the second fluctuation constrains $\bar{Q}^{\,*}_{\mathrm{int}}$ to a propensity-weighted mean of $\bar{Q}^{\,*}_{Y}$ over observed surrogates, so the cross-product $(\hat{Q}_{Y}-Q^{0}_{Y})(\hat{f}_{S}-f^{0}_{S})$ constituting $R_{SY}$ never enters the expansion.
@@ 5.3 Non-asymptotic Coverage Bound
A Berry–Esseen bound (Theorem S1, Supplement) quantifies finite-sample undercoverage of the sandwich interval in terms of bounded cluster scores, variance concentration, and the nuisance remainder, giving a minimum cluster count $J^{*}\approx 27$ for 1% maximum undercoverage at the simulation DGP. Corollary S2 shows that the constant coverage shortfall in Block I is the finite-sample signature of estimation near the product-rate boundary: the fluctuation-step variance contributes an $O(J^{-1})$ term not captured by the sandwich, and this term does not shrink relative to the leading variance term under parametric nuisance rates.
The leave-one-cluster-out jackknife for the full SA-TMLE estimator resolves this gap in practice (Section 6.3), yielding near-nominal coverage at $J\geq 30$. Formal jackknife variance consistency for targeted estimators with data-adaptive nuisance follows from the arguments of Zheng and van der Laan (2011); a complete non-asymptotic treatment is deferred to future work.
@@ 6 Monte Carlo Simulation Study
Three simulation blocks examine the main implications of the theory: stability under increasing censoring (Blocks I, III), finite-sample double-robustness under nuisance misspecification (Block II), and coverage scaling with $J$ (Block I). Throughout, “double robustness” refers to asymptotic first-order identification; finite-sample performance depends on which nuisance component is misspecified and the magnitude of the residual second-order error.
@@ 6.1 Data-Generating Process
We simulate a SW-CRT with $J$ clusters, $T=7$ calendar time steps, and $n_{j}=40$ individuals per cluster per step ($N=J\times T\times n_{j}$). Crossover times $\tau_{j}$ are drawn uniformly from $\{2,\ldots,7\}$; administrative censoring is imposed structurally so that clusters crossing at steps T1T-1 and $T$ have near-zero $g_{\Delta}$ (outcome delay $t_{\mathrm{lag}}=2$, grace period $t_{\mathrm{grace}}=1$, overall censoring rate $\approx 28\%$). The cluster random effect satisfies $b_{j}\overset{\mathrm{i.i.d.}}{\sim}\mathcal{N}(0,0.034)$ ($\mathrm{ICC}\approx 0.05$). Baseline covariates are $W_{ij,1}\sim\mathcal{N}(0,1)$, $W_{ij,2}\sim\mathrm{Bernoulli}(0.4)$, $W_{ij,3}\sim\mathrm{Uniform}(0,1)$. The secular trend is non-linear, $\lambda(t)=0.5\sin(\pi(t-1)/(T-1))+0.3((t-1)/(T-1))^{2}$, so that parametric linear-in-$t$ models are misspecified. The surrogate satisfies $S_{ijt}=0.8A_{jt}+0.4W_{ij,1}-0.3W_{ij,2}+0.5\lambda(t)+0.6b_{j}+\varepsilon^{S}_{ijt}$, $\varepsilon^{S}_{ijt}\overset{\mathrm{i.i.d.}}{\sim}\mathcal{N}(0,0.25)$; the coefficient $0.6$ on $b_{j}$ ensures $S$ is an informative bridge. The observation indicator and outcome follow
$\operatorname{logit}P(\Delta_{ijt}=1\mid S,A,W,t)$ $=3.0-2.5\cdot\mathbf{1}(t\geq T-1)-0.8\cdot\mathbf{1}(t=T)+0.4S_{ijt}+0.3A_{jt}-0.2W_{ij,3},$ (40)
$Y_{ijt}$ $=\Psi_{0}A_{jt}+0.5S_{ijt}+0.4W_{ij,1}-0.2W_{ij,2}+0.6W_{ij,3}+0.8\lambda(t)+b_{j}+\varepsilon^{Y}_{ijt},$ (41)
with $\varepsilon^{Y}_{ijt}\overset{\mathrm{i.i.d.}}{\sim}\mathcal{N}(0,0.64)$ and direct effect $\Psi_{0}=-0.28$. The true ATE is $\Psi^{*}=-0.28+0.5\times 0.8=0.12$, placed near the detection boundary for $J\approx 30$.
@@ 6.2 Simulation Scenarios
$J=30$. Three scenarios probe Theorem 2’s double-robustness: (i) outcome models misspecified (non-linear trend and $S\times A$ interaction omitted), $g_{\Delta}$ correct; (ii) $g_{\Delta}$ misspecified (surrogate $S$ omitted from the censoring model), outcome models correct; (iii) both misspecified. Scenarios (i) and (iii) are expected to show bias; scenario (ii) tests the arm of double robustness where consistent outcome models compensate for $g_{\Delta}$ misspecification.
@@ 6.3 Results
Table S5 (Supplement) reports point-estimation metrics for all three estimators across $J\in\{10,20,30,50,100\}$; the inference comparison between sandwich and jackknife for CV-TMLE is in Table 2.
@@
Point estimation. CV-TMLE achieves near-zero bias ($|\mathrm{Bias}|<0.004$) at all cluster counts, with RMSE improving steadily from 0.092 at $J=10$ to 0.030 at $J=100$. GLMM exhibits persistent bias of approximately +0.07+0.07 due to the misspecified secular time trend; its coverage collapses from 0.52 at $J=10$ to near zero at $J=100$ as the standard error shrinks around the biased point estimate. IPCW exhibits substantial positive bias (approximately +0.22+0.22) and markedly higher variance, consistent with the theoretical prediction that inverse weighting becomes unstable when $g_{\Delta}$ concentrates near zero for late-crossing clusters.
Sandwich variance and undercoverage. CV-TMLE with the sandwich interval achieves coverage of 0.87–0.89 across all $J$, well below the nominal 0.95. Table 1 diagnoses the source: the sandwich estimates $\mathrm{Var}(\mathrm{EIC}_{j})/J$ but omits the variance contribution from the fluctuation-step parameters $(\varepsilon_{Y},\varepsilon_{S})$, which vary across leave-one-cluster-out fits. The empirical variance exceeds the mean sandwich by a factor of approximately 1.6 at all $J$; this ratio is stable because both the sandwich and the omitted term are $O(J^{-1})$ under the present nuisance specification, so neither diminishes faster than the other.
Jackknife variance and coverage restoration. The leave-one-cluster-out jackknife, applied to the full two-stage estimator (re-running both fluctuation steps for each held-out cluster), captures the fluctuation-step variance that the sandwich misses. Table 1 shows that the mean jackknife variance tracks the empirical variance closely at all $J$. The resulting coverage (Table 2, Jackknife columns) is near-nominal at the practically important cluster sizes: 0.956 at $J=30$ and 0.960 at $J=50$. At very small $J$ (10 and 20) the jackknife is mildly conservative (0.970–0.980), as the leave-one-out procedure loses a larger fraction of training data per fold and slightly overestimates variability; this is a well-understood property of the jackknife in small samples and represents a safe failure mode. By $J\geq 30$ the conservatism resolves and coverage is stable near 0.956–0.968 through $J=100$.
The power tradeoff is modest: at $J=30$, jackknife power is 0.57 versus sandwich power of 0.73, reflecting wider intervals. By $J=50$, jackknife power reaches 0.786, and by $J=100$ it is 0.970 — essentially the same as the sandwich at those sizes. The undercoverage was a variance-calibration issue, not a failure of the point estimator; the cluster jackknife resolves it without altering the bias or RMSE.
@@
Table 1: Variance decomposition for CV-TMLE coverage (Block I, 500 replicates). $\widehat{\mathrm{Var}}_{\mathrm{emp}}$: empirical variance of $\hat{\Psi}$. $\overline{\mathrm{Sand.}}$: mean sandwich variance estimate. $\overline{\mathrm{Jack.}}$: mean cluster-jackknife variance estimate. Ratio $\widehat{\mathrm{Var}}_{\mathrm{emp}}/\overline{\mathrm{Sand.}}$; values >1>1 indicate sandwich underestimation attributable to unaccounted fluctuation-step variance. The jackknife variance tracks the empirical variance closely, explaining the near-nominal jackknife coverage in Table 2.
$J$ $\widehat{\mathrm{Var}}_{\mathrm{emp}}$ ($\times 10^{-3}$) $\overline{\mathrm{Sand.}}$ ($\times 10^{-3}$) $\overline{\mathrm{Jack.}}$ ($\times 10^{-3}$) Ratio Sand. Cov. Jack. Cov.
10 8.44 4.80 9.39 1.76 0.878 0.980
20 4.83 2.83 5.37 1.71 0.869 0.970
30 2.93 1.84 3.23 1.59 0.892 0.956
50 1.80 1.11 2.01 1.62 0.866 0.960
100 0.87 0.54 0.97 1.61 0.878 0.968
Table 2: Block I CV-TMLE results: point estimation and inference comparison (500 replicates, $\Psi^{*}=0.12$, $T=7$, $n_{j}=40$). Bias and RMSE are identical across both variance estimators by construction. The jackknife is the leave-one-cluster-out procedure applied to the full two-stage estimator including both fluctuation steps. Full results for GLMM and IPCW comparators in Table S5 (Supplement). MC standard error for coverage $\approx 0.010$.
Sandwich Jackknife
$J$ Bias RMSE $\widehat{\mathrm{Var}}_{\mathrm{emp}}$ ($\times 10^{-3}$) Cov. Power Cov. Power
10 +0.002+0.002 0.090 8.15 0.878 0.410 0.980 0.204
20 0.001-0.001 0.068 4.66 0.869 0.554 0.970 0.332
30 +0.003+0.003 0.054 2.91 0.892 0.727 0.956 0.570
50 +0.003+0.003 0.043 1.80 0.866 0.897 0.960 0.786
100 0.001-0.001 0.029 0.85 0.878 0.997 0.968 0.970
@@
Additional simulations varying $\sigma_{b}^{2}\in\{0.006,0.034,0.071,0.160\}$ (ICC $\approx$ 0.01–0.20) confirm that the jackknife restores near-nominal coverage (0.946–0.968) across the full empirically observed ICC range in implementation-science SW-CRTs, extending the Block I finding to the ICC heterogeneity setting; full results are in Supplement, Section S4.